find probability of each set of independent events.
Drawing a pair of black socks from a drawer of 4 brown socks, 10 black socks and 6 gray socks

Answer by ewatrrr(24785)

 
Hi,
20 socks: 4brown, 10black, 6gray
P(drawing 2 black) = 10/20*9/19 = 90/ 380 = 9/38